A normal distribution has μ = 80 and σ = 10. What is the probability of randomly selecting 75 < X < 85


1
Expert's answer
2021-06-06T15:55:16-0400

"P(75<X<85)=P(\frac{75-80}{10}<Z<{85-80}{10})=P(-0.5<Z<0.5)="

"=P(Z<0.5)-P(Z<-0.5)=P(Z<0.5)-(1-P(Z<0.5))="

"=2P(Z<0.5)-1=2*0.6915-1=0.3830."